# VerifyBench: Benchmarking Reference-based Reward Systems for Large Language Models

### v3

@@ Abstract
| Project Page | https://zju-real.github.io/VerifyBench/ | |
@@ 3.1 Construction of VerifyBench
To emulate realistic reinforcement learning (RL) scenarios involving reference-based reward systems, we curate a comprehensive collection of open-source reasoning problems paired with corresponding reference answers. These problems encompass three primary categories, general reasoning, logical reasoning and mathematical reasoning, and are aggregated from 41 distinct sources. A complete list of these data sources is provided in Appendix C.
@@
To comprehensively evaluate model performance across diverse answer formats, we define four canonical answer types: numerical values, algebraic expressions, multiple-choice selections, and free-form strings. Utilizing a general-purpose LLM Llama-3.3-70B-Instruct (Grattafiori et al., 2024), we performed automatic answer-type classification with a prompt (Appendix E.1). Questions that fall outside these categories, such as proof-based or open-ended prompts, were excluded from further analysis. Following classification, we randomly sampled 2,000 instances per answer type, resulting in a final candidate pool of 8,000 questions.
@@
We subsequently employed 22 widely used open-source and proprietary models (see Appendix D) to generate single-shot completions for the curated set of 8,000 questions, resulting in a total of 176,000 completions. To assign initial correctness labels, we utilized Llama-3.3-70B-Instruct (Grattafiori et al., 2024) within a prompt-based judgment framework. For each question, we randomly selected four completions, two labeled as correct and two labeled as incorrect by the model, and retained them for subsequent human annotation.
@@
We conducted human annotation for the aforementioned questions and their associated completions. The annotation procedure comprised two primary tasks: (1) identifying the most appropriate answer type for each question based on its formulation and corresponding ground-truth answer, and (2) evaluating the correctness of each of the four completions. Each question was independently annotated by at least two annotators. If their annotations were consistent, the labeling was finalized; otherwise, a meta-annotator resolved disagreements to ensure consistency and finalize the labels. A detailed description of the human annotation process is provided in Appendix F.
@@
Following human annotation, we identified notable biases in the models’ predictions regarding both answer types and completion correctness, leading to imbalanced data distributions. To mitigate this issue, we performed controlled downsampling to ensure uniform category-level representation and balanced correctness labels. Specifically, we retained 250 questions per answer type, resulting in a total of 1,000 questions. Each question is paired with exactly two completions, one correct and one incorrect. The resulting dataset, VerifyBench, thus comprises 2,000 well-balanced question-answer-completion-correctness tuples. Detailed statistics for VerifyBench are provided in Table 1, and illustrative examples are presented in Appendix I.
@@ 3.2 Construction of VerifyBench-Hard
To construct the dataset, we first generated single-turn completions for the queries described in Section 3.1 using a collection of 18 open-source models (Appendix D). Due to the substantial volume of generations and the associated computational costs, closed-source models were excluded from this stage. In total, we produced approximately 1.45 million completions.
@@
We subsequently subjected the selected samples to human annotation, focusing on two key aspects: identifying the answer type and determining the correctness of each completion. Each instance was annotated independently by at least two annotators. In cases where both annotators agreed, the annotation was finalized; when disagreement occurred, a meta-annotator was consulted to resolve the conflict. A detailed description of the human annotation process is provided in Appendix F.
@@ 4.1 Overall Performance
We evaluate the performance of various verification approaches on both VerifyBench and VerifyBench-Hard. For rule-based baselines, we adopt the widely used math-verify (Kydlíček, 2025) method. In the LLM-as-a-judge setting, we prompt LLMs to perform verification; detailed prompt templates are provided in Appendix E.2. Additionally, we conduct evaluations of existing model-based verifiers, employing prompts consistent with their provided templates to ensure fairness (Chen et al., 2025; Su et al., 2025; Liu et al., 2025a). Our evaluation yields several key findings:
@@ 4.2 Reference-answers Play an Important Role in Verification
The benchmark proposed in this work fundamentally differs from existing reward benchmarks by explicitly incorporating reference answers, thereby aligning more closely with the training setups of contemporary reasoning LLMs. To isolate the impact of reference answers on verification performance, we conduct an ablation study in which models are evaluated without reference answers provided in the prompt; the prompt format used is detailed in Appendix E.3.
@@ 5.1 Error Analysis
To gain deeper insights from VerifyBench, we introduce a more fine-grained taxonomy for each answer type and analyze model performance across these subcategories. This detailed analysis helps identify specific reasoning tasks or answer formats where models are particularly error-prone. We subdivide the Numeric Values category into 8 subtypes, Expressions into 5 subtypes, Multi-choice into 3 subtypes, and String into 2 subtypes. Table 5 presents the comparative performance of different models across these detailed categories. The complete taxonomy of VerifyBench and VerifyBench-Hard is provided in Appendix G.
@@
We analyzed the samples most prone to errors and identified a common underlying issue: models frequently fail to fully comprehend the question or clearly recognize the intended objective. For instance, in cases involving multi-value answers, the ordering of values is typically irrelevant. However, if the sequence of values in the model’s output differs from the golden answer, models often incorrectly classify the response as erroneous. Similarly, errors within the Expressions category, particularly involving algebraic formulas and equations, predominantly result from inadequate mathematical comprehension. Specifically, when a model outputs an unsimplified expression, superficial textual discrepancies compared to the ground-truth answer can be significant. Rather than evaluating whether the expression is mathematically equivalent upon simplification, models prematurely deem the output incorrect, thereby leading to verification failures.
@@ 5.2 Correlation Analysis
We conducted practical RL training on Qwen2.5-3B-Instruct (Qwen et al., 2025), using LLMs with varying performance on VerifyBench as LLM judgers to serve as the reward system. We then compared their improvements on downstream tasks. Specifically, we adopted Llama-3.2-3B-Instruct (Grattafiori et al., 2024), Qwen3-1.7, Qwen3-4B (Yang et al., 2025), and gpt-oss-20b (OpenAI et al., 2025) as verifiers, employed GRPO (Shao et al., 2024) as the RL algorithm, and trained on DeepMath-103K. The accuracy changes on MATH500 (Hendrycks et al., 2021), AMC23 (mathai, 2023), and MinervaMath (Lewkowycz et al., 2022) are shown in Figure 3. Meanwhile, we provide a additional correlation analysis based on Rejection Sampling Fine-tuning and all the experimental details in Appendix H.
@@ 6 Conclusion
Appendix A Limitations
@@
Appendix B Related Works
B.1 Reward Systems for Reinforcement Learning
@@
B.2 Evaluation of Reward Systems
@@
Appendix C Data Source
@@
Appendix D LLM Usage For Benchmark Construction
@@
Appendix E Prompts
E.1 Prompt for Answer Type Classification
@@ E.2 Prompt for LLM-as-a-judge
E.3 Prompt for LLM-as-a-judge without Reference
@@
Appendix F Details of Human Annotation
@@ F.1 Annotators
F.2 Instruction for Annotators
F.3 UI Design for Human Annotation Process
@@
F.3.1 Page for Primary Annotators
@@
F.3.2 Page for Meta-Annotators
@@
Appendix G Full Taxonomy of VerifyBench and VerifyBench-Hard
@@
Appendix H Correlation Experimental Details
Our correlation experiments consist of two parts: we conducted both Reinforcement Learning (RL) (Section H.1) and Rejection Sampling Fine-tuning (RFT) (Section H.2) to validate the usability of VerifyBench.
H.1 Reinforcement Learning Experiments
H.1.1 Setup
@@
For RL usage, we adopted Qwen2.5-3B-Instruct (Qwen et al., 2025) as the base model and conducted training with the DeepMath-103K (He et al., 2025) dataset. For the RL algorithm, we employed GRPO (Shao et al., 2024) optimization, setting the global batch size to 512 and the mini-batch size to 128. For each query, we sampled 16 responses to compute the group relative advantage. The maximum prompt length was set to 1024, and the maximum response length was set to 7168. For rollouts, we utilized the vLLM (Kwon et al., 2023) engine with temperature fixed at 1.0 and top_p at 1.0. Training was performed with the veRL (Sheng et al., 2025) framework using FSDP as the backend. The learning rate during training was set to 1e-6, with 20 warmup steps. For the LLM verifier, we deployed services through SGLang (Zheng et al., 2024) and performed verification via remote calls. All RL experiments were conducted on 8 H200 GPUs.
@@ H.1.2 Experimental Results
H.2 Rejection Sampling Fine-tuning Experiments
H.2.1 Setup
@@
H.2.2 Experimental Results
In our experiments, we applied rejection sampling to implement reference-based reward systems. For each question in the GSM8K and MATH training sets, we generated 64 candidate completions using Qwen2.5-Math-7B-Instruct (Yang et al., 2024) with a sampling temperature of 0.7. These responses were subsequently filtered by three verifier models with varying performance levels on VerifyBench: Llama-3.1-8B-Instruct, Qwen3-4B, and Qwen3-1.7B. Only completions consistently verified as correct were retained to form the SFT training data.
The resulting models were evaluated on multiple mathematical reasoning benchmarks, including GSM8K (Cobbe et al., 2021), MATH500 (Hendrycks et al., 2021; Lightman et al., 2023), and SVAMP (Patel et al., 2021). As shown in Figure 10, on GSM8K, MATH500, and SVAMP, using Qwen3-4B, a verifier that achieves higher accuracy on VerifyBench, consistently outperforms Llama-3.1-8B-Instruct, a lower-scoring verifier, under the same training steps. This highlights the strong alignment between VerifyBench and practical usage. Our benchmarks serve as reliable tools to guide the development of reward systems, leading to more effective training and improved model performance.
Appendix I Examples of VerifyBench
We provide some examples of VerifyBench with four different answer types in Figure 11, 12, 13 and 14. Due to space limitations, we omitted part of the text and used <omitted> as a placeholder.
Figure 11 presents a sample from the Numeric Values category, Figure 12 shows a sample from the Expression category, Figure 13 illustrates a sample from the Multi-choice category, and Figure 14 displays a sample from the String category.

### v4

@@ Abstract
| Project Page | https://zju-real.github.io/VerifyBench | |
@@ 3.1 Construction of VerifyBench
To emulate realistic reinforcement learning (RL) scenarios involving reference-based reward systems, we curate a comprehensive collection of open-source reasoning problems paired with corresponding reference answers. These problems encompass three primary categories, general reasoning, logical reasoning and mathematical reasoning, and are aggregated from 41 distinct sources. A complete list of these data sources is provided in Appendix D.
@@
To comprehensively evaluate model performance across diverse answer formats, we define four canonical answer types: numerical values, algebraic expressions, multiple-choice selections, and free-form strings. Utilizing a general-purpose LLM Llama-3.3-70B-Instruct (Grattafiori et al., 2024), we performed automatic answer-type classification with a prompt (Appendix G.1). Questions that fall outside these categories, such as proof-based or open-ended prompts, were excluded from further analysis. Following classification, we randomly sampled 2,000 instances per answer type, resulting in a final candidate pool of 8,000 questions.
@@
We subsequently employed 22 widely used open-source and proprietary models (see Appendix E) to generate single-shot completions for the curated set of 8,000 questions, resulting in a total of 176,000 completions. To assign initial correctness labels, we utilized Llama-3.3-70B-Instruct (Grattafiori et al., 2024) within a prompt-based judgment framework. For each question, we randomly selected four completions, two labeled as correct and two labeled as incorrect by the model, and retained them for subsequent human annotation.
@@
We conducted human annotation for the aforementioned questions and their associated completions. The annotation procedure comprised two primary tasks: (1) identifying the most appropriate answer type for each question based on its formulation and corresponding ground-truth answer, and (2) evaluating the correctness of each of the four completions. Each question was independently annotated by at least two annotators. If their annotations were consistent, the labeling was finalized; otherwise, a meta-annotator resolved disagreements to ensure consistency and finalize the labels. A detailed description of the human annotation process is provided in Appendix H .
@@
Following human annotation, we identified notable biases in the models’ predictions regarding both answer types and completion correctness, leading to imbalanced data distributions. To mitigate this issue, we performed controlled downsampling to ensure uniform category-level representation and balanced correctness labels. Specifically, we retained 250 questions per answer type, resulting in a total of 1,000 questions. Each question is paired with exactly two completions, one correct and one incorrect, aimed to minimize bias arising from class imbalance. In this way, we maintain an equal number of samples across these 4 answer types in VerifyBench, enabling us to assess a model’s verification accuracy under different scenarios. The resulting dataset, VerifyBench, thus comprises 2,000 well-balanced question-answer-completion-correctness tuples. Detailed statistics for VerifyBench are provided in Table 1 and Appendix F, and illustrative examples are presented in Appendix K and L.
@@ 3.2 Construction of VerifyBench-Hard
To construct the dataset, we first generated single-turn completions for the queries described in Section 3.1 using a collection of 18 open-source models (Appendix E). Due to the substantial volume of generations and the associated computational costs, closed-source models were excluded from this stage. In total, we produced approximately 1.45 million completions.
@@
We subsequently subjected the selected samples to human annotation, focusing on two key aspects: identifying the answer type and determining the correctness of each completion. Each instance was annotated independently by at least two annotators. In cases where both annotators agreed, the annotation was finalized; when disagreement occurred, a meta-annotator was consulted to resolve the conflict. A detailed description of the human annotation process is provided in Appendix H.
@@ 4.1 Overall Performance
We evaluate the performance of various verification approaches on both VerifyBench and VerifyBench-Hard. For rule-based baselines, we adopt the widely used math-verify (Kydlíček, 2025) method111For math-verify, we report the accuracy of version v0.8.0.. In the LLM-as-a-judge setting, we prompt general LLMs to perform verification; detailed prompt templates are provided in Appendix G.2. Additionally, we conduct evaluations of existing verification-specialized model-based verifiers, which have been carefully trained on verification tasks, employing prompts consistent with their provided templates to ensure fairness (Chen et al., 2025; Su et al., 2025; Liu et al., 2025a). Our evaluation yields several key findings:
@@ 4.2 Reference-answers Play an Important Role in Verification
The benchmark proposed in this work fundamentally differs from existing reward benchmarks by explicitly incorporating reference answers, thereby aligning more closely with the training setups of contemporary reasoning LLMs. To isolate the impact of reference answers on verification performance, we conduct an ablation study in which models are evaluated without reference answers provided in the prompt; the prompt format used is detailed in Appendix G.3.
@@ 5.1 Error Analysis
To gain deeper insights from VerifyBench, we introduce a more fine-grained taxonomy for each answer type and analyze model performance across these subcategories. This detailed analysis helps identify specific reasoning tasks or answer formats where models are particularly error-prone. We subdivide the Numeric Values category into 8 subtypes, Expressions into 5 subtypes, Multi-choice into 3 subtypes, and String into 2 subtypes. Table 5 presents the comparative performance of different models across these detailed categories. The complete taxonomy of VerifyBench and VerifyBench-Hard is provided in Appendix I.
@@
We analyzed the samples most prone to errors and identified a common underlying issue: models frequently fail to fully comprehend the question or clearly recognize the intended objective. For instance, in cases involving multi-value answers, the ordering of values is typically irrelevant. However, if the sequence of values in the model’s output differs from the golden answer, models often incorrectly classify the response as erroneous. Similarly, errors within the Expressions category, particularly involving algebraic formulas and equations, predominantly result from inadequate mathematical comprehension. Specifically, when a model outputs an unsimplified expression, superficial textual discrepancies compared to the ground-truth answer can be significant. Rather than evaluating whether the expression is mathematically equivalent upon simplification, models prematurely deem the output incorrect, thereby leading to verification failures. We provide, in Appendix M, examples of misjudgments made by gpt-oss-120b, the best-performing open-source model on VerifyBench within each category. These examples span different answer types and are intended to help better understand why LLM verifiers make errors.
@@ 5.2 Correlation Analysis
We conducted practical RL training on Qwen2.5-3B-Instruct (Qwen et al., 2025), using LLMs with varying performance on VerifyBench as LLM judgers to serve as the reward system. We then compared their improvements on downstream tasks. Specifically, we adopted math-verify (Kydlíček, 2025), Llama-3.2-3B-Instruct (Grattafiori et al., 2024), Qwen3-1.7, Qwen3-4B (Yang et al., 2025), and gpt-oss-20b (OpenAI et al., 2025) as verifiers, employed GRPO (Shao et al., 2024) as the RL algorithm, and trained on DeepMath-103K. The accuracy changes on MATH500 (Hendrycks et al., 2021), AMC23 (mathai, 2023), and MinervaMath (Lewkowycz et al., 2022) are shown in Figure 3. Meanwhile, we provide a additional correlation analysis based on Rejection Sampling Fine-tuning and all the experimental details in Appendix J.
@@ 6 Conclusion
Acknowledgement
This work was supported by National Natural Science Foundation of China (No. 62506332), “Pioneer” and “Leading Goose” R&D Program of Zhejiang (NO. 2026C02A1223) and CCF-Tencent Rhino-Bird Open Research Fund.
@@ Acknowledgement
Reproducibility statement
We provide the VerifyBench dataset and evaluation code in the supplementary materials to facilitate the research community’s use of our benchmark. In Appendix J, we detail the hyperparameter settings used in our evaluations, enabling researchers to reproduce our results. For the RL experiments in Section 5.2, we conducted experiments using the ‘remote_rm‘ recipe provided by veRL (Sheng et al., 2025). The detailed experimental configurations are also included in Appendix J to support reproducibility. Nevertheless, due to the inherent stochasticity in LLM training and inference, results may still exhibit variability.
Appendix A LLM Usage Declaration
In writing this paper, we only used LLMs for polishing. The generation of ideas in this work did not involve any assistance from LLMs. The experimental design and manuscript writing were not directly produced by LLMs either. The models were used solely as a polishing tool: specifically, we first drafted the manuscript, then refined it with the help of an LLM, and finally the authors conducted another round of verification after polishing.
Appendix B Limitations
@@
Appendix C Related Works
C.1 Reward Systems for Reinforcement Learning
@@
C.2 Evaluation of Reward Systems
@@
Appendix D Data Source
@@
Appendix E LLM Usage For Benchmark Construction
@@
Appendix F Detailed Benchmark Statistics
In this section, we present additional statistics for VerifyBench and VerifyBench-Hard, including the distribution of completion source models for both benchmarks (Appendix F.1) and the distribution of token lengths (Appendix F.2).
F.1 Source Model Distribution
VerifyBench and VerifyBench-Hard include completions generated by different source models. To better understand their distribution, we report the distribution of source models in both benchmarks, as shown in Table 8.
| Series | Model | VerifyBench | VerifyBench-Hard | ||
| # of samples | ratio (%) | # of samples | ratio (%) | ||
| OpenAI | gpt-4o-2024-11-20 | 91 | 4.55 | 0 | 0.00 |
| gpt-4o-mini | 79 | 3.95 | 0 | 0.00 | |
| anthropic | claude-3.7-sonnet | 50 | 2.50 | 0 | 0.00 |
| deepseek-math | deepseek-math-7b-instruct | 107 | 5.35 | 101 | 10.10 |
| deepseek-math-7b-rl | 92 | 4.60 | 72 | 7.20 | |
| DeepSeek | DeepSeek-V3 | 80 | 4.00 | 0 | 0.00 |
| DeepSeek-R1 | 101 | 5.05 | 0 | 0.00 | |
| DeepSeek-R1-Distill-Qwen-7B | 85 | 4.25 | 0 | 0.00 | |
| DeepSeek-R1-Distill-Qwen-32B | 92 | 4.60 | 112 | 11.20 | |
| gemma-3 | gemma-3-1b-it | 148 | 7.40 | 123 | 12.30 |
| gemma-3-4b-it | 87 | 4.35 | 80 | 8.00 | |
| gemma-3-12b-it | 87 | 4.35 | 7 | 0.70 | |
| Llama-3 | Llama-3.3-70B-Instruct | 84 | 4.20 | 0 | 0.00 |
| Llama-3-8B-Instruct | 117 | 5.85 | 56 | 5.60 | |
| Qwen2.5 | Qwen2.5-7B-Instruct | 82 | 4.10 | 106 | 10.60 |
| Qwen2.5-72B-Instruct | 86 | 4.30 | 18 | 1.80 | |
| Qwen2.5-Math | Qwen2.5-Math-1.5B-Instruct | 86 | 4.30 | 49 | 4.90 |
| Qwen2.5-Math-7B-Instruct | 83 | 4.15 | 91 | 9.10 | |
| Qwen2.5-Math-72B-Instruct | 77 | 3.85 | 21 | 2.10 | |
| QwQ | QwQ-32B | 120 | 6.00 | 50 | 5.00 |
| Yi-1.5 | Yi-1.5-9B-Chat-16K | 92 | 4.60 | 21 | 2.10 |
| Yi-1.5-34B-Chat | 74 | 3.70 | 93 | 9.30 | |
On VerifyBench and VerifyBench-Hard, the distribution of completions produced by different models does not exhibit extreme imbalance (each model contributes roughly 2–7% on VerifyBench, and 1-12% on VerifyBench-Hard).
F.2 Token Length Distribution of Completions in VerifyBench(-Hard)
We provide the token-length distributions of completions in both VerifyBench and VerifyBench-Hard, as shown in Figure 4.
It can be seen that the token-length range of the completions in our constructed VerifyBench is broad: it includes shorter chat-style completions as well as longer R1-like completions. This diversity improves the coverage and realism of VerifyBench, making it a more reliable benchmark.
Appendix G Prompt Templates
G.1 Prompt Template for Answer Type Classification
@@ G.2 Prompt Template for LLM-as-a-judge
G.3 Prompt Template for LLM-as-a-judge without Reference
@@
Appendix H Details of Human Annotation
@@ H.2 Annotators
H.3 UI Design for Human Annotation Process
@@
H.3.1 Page for Primary Annotators
@@
H.3.2 Page for Meta-Annotators
@@
Appendix I Full Taxonomy of VerifyBench and VerifyBench-Hard
@@
Appendix J Correlation Experimental Details
Our correlation experiments consist of two parts: we conducted both Reinforcement Learning (RL) (Section J.1) and Rejection Sampling Fine-tuning (RFT) (Section J.2) to validate the usability of VerifyBench.
J.1 Reinforcement Learning Experiments
J.1.1 Setup
@@
For RL experiments, we adopted Qwen2.5-3B-Instruct (Qwen et al., 2025) as the base model and conducted training with the DeepMath-103K (He et al., 2025) dataset. For the RL algorithm, we employed GRPO (Shao et al., 2024) optimization, setting the global batch size to 512 and the mini-batch size to 128. For each query, we sampled 16 responses to compute the group relative advantage. The maximum prompt length was set to 1024, and the maximum response length was set to 7168. For rollouts, we utilized the vLLM (Kwon et al., 2023) engine with temperature fixed at 1.0 and top_p at 1.0. Training was performed with the veRL (Sheng et al., 2025) framework using FSDP as the backend. The learning rate during training was set to 1e-6, with 20 warmup steps. For the LLM verifier, we deployed services through SGLang (Zheng et al., 2024) and performed verification via remote calls. All RL experiments were conducted on 8 H200 GPUs.
@@
J.2 Rejection Sampling Fine-tuning Experiments
J.2.1 Setup
@@
J.2.2 Experimental Results
In our experiments, we applied rejection sampling to implement reference-based reward systems. For each question in the GSM8K and MATH training sets, we generated 64 candidate completions using Qwen2.5-Math-7B-Instruct (Yang et al., 2024) with a sampling temperature of 0.7. These responses were subsequently filtered by 4 verifier models with varying performance levels, on VerifyBench: Llama-3.1-8B-Instruct, Qwen3-4B, Qwen3-1.7B and gpt-oss-20b, and a rule-based verifier, math-verify. Only completions consistently verified as correct were retained to form the SFT training data.
The resulting models were evaluated on multiple mathematical reasoning benchmarks, including GSM8K (Cobbe et al., 2021), MATH500 (Hendrycks et al., 2021; Lightman et al., 2023), and SVAMP (Patel et al., 2021). As shown in Figure 11, on GSM8K, MATH500, and SVAMP, using gpt-oss-20b, a verifier that achieves higher accuracy on VerifyBench, consistently outperforms Llama-3.1-8B-Instruct, a lower-scoring verifier, under the same training steps. This highlights the strong alignment between VerifyBench and practical usage. Our benchmarks serve as reliable tools to guide the development of reward systems, leading to more effective training and improved model performance.
Appendix K Examples of VerifyBench
In this section, we provide representative samples from the four answer-type categories in VerifyBench to facilitate a clearer understanding of the benchmark we constructed. Due to space limitations, we omitted part of the text and used <omitted> as a placeholder.
Figure K.1 presents a sample from the Numeric Values category, Figure K.2 shows a sample from the Expression category, Figure K.3 illustrates a sample from the Multi-choice category, and Figure K.4 displays a sample from the String category.
K.1 An example of VerifyBench (Answer type: Numeric Values)
K.2 An example of VerifyBench (Answer type: Expression)
K.3 An example of VerifyBench (Answer type: Multi-choice)
K.4 An example of VerifyBench (Answer type: String)
Appendix L Examples of VerifyBench-Hard
In this section, we present hard-to-verify samples from the four different answer-type categories in VerifyBench-Hard, enabling a clearer understanding of the construction and characteristics of VerifyBench-Hard. Due to space limitations, we omitted part of the text and used <omitted> as a placeholder.
Figure L.1 presents a sample from the Numeric Values category, Figure L.2 shows a sample from the Expression category, Figure L.3 illustrates a sample from the Multi-choice category, and Figure L.4 displays a sample from the String category.
L.1 An example of VerifyBench-Hard (Answer type: Numeric Values)
L.2 An example of VerifyBench-Hard (Answer type: Expression)
L.3 An example of VerifyBench-Hard (Answer type: Multi-choice)
L.4 An example of VerifyBench-Hard (Answer type: String)
Appendix M Error Case Study
To facilitate a deeper understanding of the error patterns exhibited by current models on VerifyBench, we present four representative failure cases of gpt-oss-20b. These examples span multiple answer types, illustrating the breadth of scenarios in which the model’s verification judgments break down.